# Here Be Dragons: Bimodal posteriors arise from numerical integration error in longitudinal models

### v2

@@ Abstract
Longitudinal models with dynamics governed by differential equations may require numerical integration alongside parameter estimation. We have identified a situation where the numerical integration introduces error in such a way that it becomes a novel source of non-uniqueness in estimation. We obtain two very different sets of parameters, one of which is a good estimate of the true values and the other a very poor one. The two estimates have forward numerical projections statistically indistinguishable from each other because of numerical error. In such cases, the posterior distribution for parameters is bimodal, with a dominant mode closer to the true parameter value, and a second cluster around the errant value. We demonstrate that multi-modality exists both theoretically and empirically for an affine first order differential equation, that a simulation workflow can test for evidence of the issue more generally, and that Markov Chain Monte Carlo sampling with a suitable solution can avoid bimodality. The issue of multi-modal posteriors arising from numerical error has consequences for Bayesian inverse methods that rely on numerical integration more broadly.
@@ 1 Introduction
The use of numerical integration in methods for estimating differential equation parameters from observations introduces a source of error (Butcher,, 2016) that is distinct from other processes such as measurement error (Agapiou et al.,, 2014). In the Bayesian literature, the effect of numerical error on the posterior distribution is under-explored and we wish to present a case study on identifiability problems which doubles as a procedure for investigating them under similar circumstances. In this paper, we describe a specific pathology arising from numerical error: erroneous multi-modality of posterior parameter distributions for affine first order ODEs where there are multiple parameter combinations that produce the same numerical solution. Such multi-modality is a previously undescribed form of non-identifiability in Bayesian methods.
@@
We first observed bimodality arising from numerical methods while testing a longitudinal model for the von Bertalanffy function (Von Bertalanffy,, 1938). Posterior bimodality leads to independent Markov Chain Monte Carlo (MCMC) chains converging to very different, but very specific, parameter combinations, one which is a good estimate of the true values (i.e. parameters known from simulation), and a separate very bad estimate. For the erroneous combination, under the chosen numerical method and step size, forward projection produces estimates of $Y(t_{j})$ that are the same as those produced by the correct parameters, due to numerical error. The posterior is, after all, conditioned on the numerical method used in the model, not the true solution to the integral. As our case study arose from ecological literature, We are thinking about growth functions for DEs, and growth parameters as what is being estimated. However the example is a simple DE likely to appear in other examples. The language we use will reflect the ecological background but the situation is more general.
@@
Non-identifiability is a general statistical concept for a situation where multiple estimate values have an indistinguishable relationship to data (Rothenberg,, 1971). The form of non-identifiability We are dealing with is different to cases described in the literature. For example Gelfand and Sahu, (1999) discusses situations where the set of observations is small relative to the parameter space such as can arise in generalised linear mixed effects models. Auger-Méthé et al., (2016) addresses large measurement error as a potential source, and Cockayne et al., (2019) discusses a situation where the posterior distribution is bimodal because there are two true solutions to the underlying function. Numerical integration error as a source of estimation problems is rarely addressed in the existing literature. The only mentions we found were Agapiou et al., (2014) looking at convergence for MCMC at different levels of discretisation and Cotter et al., (2010) where the step size of an Euler method was found to bias estimates. In supplementary material to O’Brien et al., (2024) we investigated bias arising from numerical error for a linear ODE. Posterior multi-modality arising from numerical methods appears to be completely novel.
@@
When multi-modality arises from numerical error, stability of the integration method matters for potentially a very large region of the parameter space – far beyond true or even plausible parameter combinations – the boundaries of which are probably unknown when first going in. The numerical methods we use (Runge-Kutta 4 and 4-5 order) are extremely common (e.g. Falster et al.,, 2017), and numerically more reliable than other commonly used methods such as Euler (implemented in Cotter et al.,, 2010 and Iida et al.,, 2014). We demonstrate that the problem arises theoretically, and is empirically observable, for an affine first-order ordinary differential equation with negative slope, such as the von Bertalanffy (Von Bertalanffy,, 1938) model where we first observed it. The von Bertalanffy model is our central example in a re-parameterised form. Affine ODEs are common, and the RK4 method is typically considered reliable for short step sizes (Butcher,, 2016). To find a situation where the combination of the two produced a novel pathology was unexpected. The affine ODE has an analytic solution which allows usto side-step the identifiability problem by not using numerical methods, however if numerical error introduces non-uniqueness elsewhere it may be in situations without such a work-around.
@@
In this paper we first prove that in general, affine ODEs subject to RK4 numerical integration in a longitudinal model will have multi-modal posteriors. We then present a simulation workflow for matching empirical results that allows a user to test the parameter space of more general models for multi-modal pathologies before embarking on large-scale analysis. Our testing starts with MCMC estimation, with different numerical methods, priors, and parameter combinations which still produce bimodal posteriors. The problem persists for deterministic estimation methods that use gradient ascent to maximise the likelihood, such as the L-BFGS algorithm (Nocedal and Wright,, 2006). We report the results of experiments to demonstrate how the problem appears in practice, and investigate whether common approaches to dealing with such issues work. We have another function of interest – the extremely non-linear Canham function based on a log-normal distribution function from Canham et al., (2004) – so we use the simulation workflow to test whether we find evidence of bimodal posteriors for parameters in that case as well.
@@
Our use of simulation is distinct from a simulation-based calibration method (Talts et al.,, 2020; Modrák et al.,, 2023) which uses data simulated from the chosen prior to test the consistency of the posterior with rank statistics. We are doing a predominantly qualitative rather than quantitative analysis testing for evidence of the existence of posterior bimodality arising from numerical error, and use simulated data that does not come from the chosen priors. We consider this a reasonable approach because our intent is to demonstrate the existence and mechanism by which numerical bimodality arises, we already have a fix in the affine ODE’s analytic solution. In the spirit of Gelman et al., (2020), deliberately breaking a model can be informative, and we aim to present a generally applicable workflow using simulation for testing whether an intended model is susceptible to these novel problems.
@@ 2 Theoretical analysis
We demonstrate the bimodal posterior estimate behaviour for dynamics based on the affine first-order ordinary differential equation (ODE)
$f(Y(t),\beta_{0},\beta_{1})=\beta_{0}-\beta_{1}Y(t),$ (2.1)
@@
[…]
Equation (2.1) is considered a particularly pathological example for numerical methods (Butcher,, 2016) as it is unbounded above and below, and the negative slope in particular means that negative growth increments can arise from the numerics even if the underlying system is strictly positive. The move from continuous time in the analytic solution to discrete time in the numerical one can result in estimates of $Y(t)$ oscillating around the asymptotic size as the calculated gradient overshoots 0 then is pulled back by being a negative value as seen in Figure 1. Equation (2.1) serves as an ideal example of the bimodal posterior because of the known pathologies, it is simple with two parameters, and has an analytic solution which allows us to easily access true values for $Y(t)$ without introducing compounding numerical error. The analytic solution is
$Y(t)=\frac{\beta_{0}}{\beta_{1}}+\bigg{(}Y(0)-\frac{\beta_{0}}{\beta_{1}}\bigg{)}\exp(-\beta_{1}t),$ (2.3)
@@
We choose the initial condition $0<Y(0)=1<\frac{\beta_{0}}{\beta_{1}}$. Equation (2.3) gives a mathematical model that is defined for regions that may not correspond to realistic behaviour in the system being modelled. For example, if $Y$ represents physical size, $Y(t)<0$ is impossible. Negative growth is also possible if $Y(t)>\frac{\beta_{0}}{\beta_{1}}$.
Figure 1: Behaviour of numerical solutions at erroneous and true parameter locations showing considerable instability at the erroneous point. The true $Y(t_{j})$ values are those of the analytic solution with $\beta_{0}=10$ and $\beta_{1}=1$ that the model attempts to estimate. (a) and (b) compare the estimated sizes for integer times between 0 and 4 to the analytic solution to show both the good and bad parameters give numerical values very close to the analytic ones. In (c) and (d) we see the sub-steps as well, which are connected up based on the numerical step they come from. The bad parameter estimates demonstrate considerable instability and even a negative value. (e) and (f) give the gradients at the numerical sub-steps and show why the estimated sizes are so unstable for the bad estimate as the gradient oscillates about 0.
@@
[…]
We are attempting to estimate the parameters $\beta_{0}$ and $\beta_{1}$ from simulated observations $y_{j}$ of the form in Equation (2.4). We estimate $\hat{Y}_{j}$ given the prior distribution
@@
and the longitudinal model in Equation (2.2), which encodes the auto-correlation in the data. For this demonstration We are looking at a single simulated individual. To make fitting the model easier we apply a translation by the mean observed size $\bar{y}$, which produces the implemented DE
@@
so the output estimates are consistent with Equation (2.1). We have default prior distributions for the parameters which are
@@
where the mean and standard deviation for $\beta_{1}$ are for the underlying log-transformed normal distribution. Sensitivity of estimation to the standard deviation is tested as part of the experiments. The minimal hierarchical set-up is based on a single individual structure such as used in O’Brien et al., (2024), but with a number chosen for the error standard deviation instead of fitting an error parameter.
@@ 2.1 Numerical integration theory
In this section we prove non-identifiability of the affine model parameters under a Runge-Kutta 4th order numerical method (RK4) in a theoretical context using the true values rather than observed data. We prove that, amongst the set of solutions for which the asymptotic size is $\alpha$, the affine ODE has multiple $\hat{\beta}_{1}$ parameters that will produce the same sizes over time for the RK4 method, then look at the specific case study where $\beta_{0}=10$ and $\beta_{1}=1$.
@@
In the case of Equation (2.1) the time parameter $t$ does not appear explicitly so our $k$ formulas do not depend on it. We will refer to where the $k$ parameters are estimated as ‘sub-steps’. Note that $t_{j}$ and $t_{j+1}=t_{j}+h$ are typically different to the times at which observations occur. The observation interval is an upper bound on the numerical step size $h$, but we can take multiple numerical steps between the observations.
In a chosen situation where $\beta_{0}=10$ and $\beta_{1}=1$ which produces a bimodal posterior, at the second (erroneous) parameter combination the numerical method is sufficiently unstable that forward projection produces $\hat{Y}(t_{j})$ values that match the analytic solution at the true parameters. For a visual comparison see Figure 1(a) and (b), which leverages the bad estimates ($\beta_{0}=49.3,\beta_{1}=4.91$) paired to our chosen parameters $\beta_{0}=10$ and $\beta_{1}=1$. The estimation algorithm will converge to any parameter combination that gives $\hat{Y}(t_{j})$ close to the observations provided. In these cases the algorithm finds a point far from the true value where size estimates match the observations extremely well because of numerical integration error. Non-identifiability here is not a problem with MCMC or another chosen estimation method, it is a structural issue arising from the use of numerical integration in the longitudinal structure.
@@
The following provides proofs for non-identifiability in general for affine ODEs when using the RK4 numerical method. In the case of the affine ODE we can leverage an explicit from of the RK4 algorithm in Equation (2.5) to get a polynomial in $\beta_{0}$ and $\beta_{1}$. We prove non-identifiability for solutions that satisfy:
@@
For a first order affine ODE of the form
@@
$\hat{Y}(t_{j}+h)=Y(t_{j})+(\alpha-Y(t_{j}))\bigg{(}h\beta_{1}-\frac{h^{2}}{2}\beta_{1}^{2}+\frac{h^{3}}{6}\beta_{1}^{3}-\frac{h^{4}}{24}\beta_{1}^{4}\bigg{)}$ (2.6)
@@
First we calculate formulas for each $k_{i}$ that depend only on $\beta$s and $Y(t_{j})$ by substituting in the function definition sequentially. We start with
@@
To extract the polynomial we substitute the $k$s into Equation 2.7 and do some re-organising of terms to get
@@
The affine DE with negative slope always converges to the asymptotic size $\beta_{0}/\beta_{1}=\alpha$, and in fact if we make the substitution $\alpha\beta_{1}=\beta_{0}$ we can reduce to a function of a single variable $\beta_{1}$ to get
@@
For an affine first order ODE of the form
@@
$10=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-$ 1 . (2.8)
@@
From the analytic solution in Equation (2.3) we know that
@@
which we can substitute into Equation (2.9) to get
@@
$=$ $\ (Y(t_{j})-\alpha)(e^{-\beta_{1}h}-1).$ (2.10)
When solving for $\hat{\beta}_{1}$ we re-arrange Equation (2.10) to get
@@
$=$ $\ (\alpha-Y(t_{j}))\big{(}f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1\big{)}.$ (2.11)
The only variable in Equation (2.11) is $\hat{\beta}_{1}$, so the solutions will be the same as those that solve
$f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1=0.$ (2.12)
@@
$Y(t_{j}+nh)-Y(t_{j})=f(\hat{\beta}_{1},h)\Bigg{(}\sum_{i=0}^{n-1}\alpha-Y(t_{j}+ih)\Bigg{)}.$ (2.13)
As in the 1-step case we take the left hand side (LHS) of Equation (2.13) and substitute in the analytic solution $Y(t_{j}+nh)=\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}nh}$ to get
$\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}nh}-Y(t_{j})=(Y(t_{j})-\alpha)(e^{-\beta_{1}nh}-1),$ (2.14)
@@
$e^{-\beta_{1}nh}-1=\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(e^{-\beta_{1}h}-1),$ (2.15)
@@
Substituting Equation (2.15) into Equation (2.14) gives
$\text{LHS}=(Y(t_{j})-\alpha)\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(e^{-\beta_{1}h}-1).$ (2.16)
Let us return to the right side of Equation (2.13), and look at the argument within the sum. We substitute in the analytic solution to give
@@
$\text{RHS}=f(\hat{\beta}_{1},h)(\alpha-Y(t_{j}))\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}.$ (2.17)
Recombining Equations (2.16) and (2.17), we subtract Equation (2.16) from both sides to give
@@
$=$ $\ (\alpha-Y(t_{j}))\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1).$ (2.18)
Equation (2.18) is 0 when
$0=f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-1,$ (2.19)
which is the same as Equation (2.12) from the 1-step case. As such, $\hat{\beta}_{1}$ solutions will be the same for all $n$. Equation (2.19) does not depend on $t_{j}$, so as long as $Y(t_{j})\neq\alpha$ the solutions are independent of $Y(t_{j})$ as well. ∎
@@
For an affine first order ODE of the form
@@
$10=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-$ 1
@@
For some parameter values Equation (2.19) may have a combination of real and complex roots, but as there must be at least one solution which approximates with the true (real) parameter $\beta_{1}$, we can exclude the case of all imaginary roots. If the discriminant of the polynomial in $\hat{\beta}_{1}$ is 0 there are some special cases where one real root (of some multiplicity) exists, that can be determined with further calculations when $h$ and $\beta_{1}$ are known. For example, for a given $h$ there will be one $\beta_{1}$ such that there is a single real root of multiplicity four given by
@@
and $\beta_{1}$ can be determined through calculation using the discriminant when $h$ is chosen (Rees01021922). ∎
@@
For an affine first order ODE of the form
@@
As the true value $\hat{\beta_{1}}=\beta_{1}$ does not perfectly match the numeric solution to the analytic one, the estimate closest to $\beta_{1}$ that does solve Equation (2.19) will not be $\hat{\beta_{1}}=\beta_{1}$. ∎

### v3

@@ Abstract
Longitudinal models with dynamics governed by differential equations may require numerical integration alongside parameter estimation. We have identified a situation where the numerical integration introduces error in such a way that it becomes a novel source of non-uniqueness in estimation. We obtain two very different sets of parameters, one of which is a good estimate of the true values and the other a very poor one. The two estimates have forward numerical projections statistically indistinguishable from each other because of numerical error. In such cases, the posterior distribution for parameters is bimodal, with a dominant mode closer to the true parameter value, and a second cluster around the errant value. We demonstrate that multi-modality exists both theoretically and empirically for a linear first order differential equation, that a simulation workflow can test for evidence of the issue more generally, and that Markov Chain Monte Carlo sampling with a suitable solution can avoid bimodality. The issue of multi-modal posteriors arising from numerical error has consequences for Bayesian inverse methods that rely on numerical integration more broadly.
@@ 1 Introduction
The use of numerical integration in methods for estimating differential equation parameters from observations introduces a source of error (Butcher,, 2016) that is distinct from other processes such as measurement error (Agapiou et al.,, 2014). In the Bayesian literature, the effect of numerical error on the posterior distribution is under-explored and we wish to present a case study on identifiability problems which doubles as a procedure for investigating them under similar circumstances. In this paper, we describe a specific pathology arising from numerical error: erroneous multi-modality of posterior parameter distributions for linear first order ODEs where there are multiple parameter combinations that produce the same numerical solution. Such multi-modality is a previously undescribed form of non-identifiability in Bayesian methods.
@@
We first observed bimodality arising from numerical methods while testing a longitudinal model for the von Bertalanffy function (Von Bertalanffy,, 1938). Posterior bimodality leads to independent Markov Chain Monte Carlo (MCMC) chains converging to very different, but very specific, parameter combinations, one which is a good estimate of the true values (i.e. parameters known from simulation), and a separate very bad estimate. For the erroneous combination, under the chosen numerical method and step size, forward projection produces estimates of $Y(t_{j})$ that are the same as those produced by the correct parameters, due to numerical error. The posterior is, after all, conditioned on the numerical method used in the model, not the true solution to the integral. In general, where numerical methods are implemented in estimation they should be treated as part of the statistical model. As our case study arose from ecological literature, we are thinking about growth functions for DEs, and growth parameters as what is being estimated. However the example is a simple DE likely to appear in other examples. The language we use will reflect the ecological background but the situation is more general.
@@
Non-identifiability is a general statistical concept for a situation where multiple estimate values have an indistinguishable relationship to data (Rothenberg,, 1971). The form of non-identifiability we are dealing with is different to cases described in the literature. For example Gelfand and Sahu, (1999) discusses situations where the set of observations is small relative to the parameter space such as can arise in generalised linear mixed effects models. Auger-Méthé et al., (2016) addresses large measurement error as a potential source, and Cockayne et al., (2019) discusses a situation where the posterior distribution is bimodal because there are two true solutions to the underlying function. Numerical integration error as a source of estimation problems is rarely addressed in the existing literature. The only mentions we found were Agapiou et al., (2014) looking at convergence for MCMC at different levels of discretisation and Cotter et al., (2010) where the step size of an Euler method was found to bias estimates. Supplementary material to O’Brien et al., (2024) investigated bias arising from numerical error for a linear ODE. Posterior multi-modality arising from numerical error appears to be completely novel.
@@
When multi-modality arises from numerical error, stability of the integration method matters for potentially a very large region of the parameter space – far beyond true or even plausible parameter combinations – the boundaries of which are probably unknown when first going in. The numerical methods we use (Runge-Kutta 4 and 4-5 order) are extremely common (e.g. Falster et al.,, 2017), and numerically more reliable than other commonly used methods such as Euler (implemented in Cotter et al.,, 2010 and Iida et al.,, 2014). We demonstrate that the problem arises theoretically, and is empirically observable, for a linear first-order ordinary differential equation with negative slope, such as the von Bertalanffy (Von Bertalanffy,, 1938) model where we first observed it. The von Bertalanffy model is our central example in a re-parameterised form. Linear ODEs are common, and the RK4 method is typically considered reliable for short step sizes (Butcher,, 2016). To find a situation where the combination of the two produced a novel pathology was unexpected. The linear ODE has an analytic solution which allows me to side-step the identifiability problem by not using numerical methods, however if numerical error introduces non-uniqueness elsewhere it may be in situations without such a work-around.
@@
In this paper we first prove that in general, linear first order ODEs subject to RK4 numerical integration in a longitudinal model will have multi-modal posteriors. We then present a simulation workflow for matching empirical results that allows a user to test the parameter space of more general models for multi-modal pathologies before embarking on large-scale analysis. Our testing uses MCMC estimation with different true parameter values. Supplements A and B show that different prior parameters and estimation methods do not change the problem. We report the results of experiments to demonstrate how the problem appears in practice, and investigate whether common approaches to dealing with such issues work. We have another function of interest – the extremely non-linear Canham function based on a log-normal distribution function from Canham et al., (2004) – and in Supplement C: Canham Testing we use the simulation workflow to test whether there is evidence of bimodal posteriors for parameters in that case as well.
@@
Our use of simulation is distinct from a simulation-based calibration method (Talts et al.,, 2020; Modrák et al.,, 2023) which uses data simulated from the chosen prior to test the consistency of the posterior with rank statistics. We are doing a predominantly qualitative rather than quantitative analysis testing for evidence of the existence of posterior bimodality arising from numerical error, and use simulated data that does not come from the chosen priors. We consider this a reasonable approach because our intent is to demonstrate the existence and mechanism by which numerical bimodality arises, we already have a fix in the linear ODE’s analytic solution. In the spirit of Gelman et al., (2020), deliberately breaking a model can be informative, and we aim to present a generally applicable workflow using simulation for testing whether an intended model is susceptible to these novel problems.
@@ 2 Theoretical analysis
We demonstrate the bimodal posterior estimate behaviour for dynamics based on the linear first-order ordinary differential equation (ODE)
$\frac{dY}{dt}=f(Y(t),\beta_{0},\beta_{1})=\beta_{0}-\beta_{1}Y(t),$ (1)
@@
$Y(t_{j+1})=Y(t_{j})+\int_{t_{j}}^{t_{j+1}}f(Y(t),\beta_{0},\beta_{1})\,dt.$ (2)
Equation (1) is considered a particularly pathological example for numerical methods (Butcher,, 2016) as it is unbounded above and below, and the negative slope in particular means that negative growth increments can arise from the numerics even if the underlying system is strictly positive. The move from continuous time in the analytic solution to discrete time in the numerical one can result in estimates of $Y(t)$ oscillating around the asymptotic size. The oscillation arises when the calculated gradient overshoots 0 then is pulled back as seen in Figure 1. Equation (1) serves as an ideal example of the bimodal posterior because of the known pathologies, it is simple with two parameters, and has an analytic solution which allows us to easily access true values for $Y(t)$ without introducing compounding numerical error. The known analytic solution for a linear model is
$Y(t)=\frac{\beta_{0}}{\beta_{1}}+\bigg(Y(0)-\frac{\beta_{0}}{\beta_{1}}\bigg)\exp(-\beta_{1}t),$ (3)
@@
We choose the initial condition $0<Y(0)=1<\frac{\beta_{0}}{\beta_{1}}$. Equation (3) gives a mathematical model that is defined for regions that may not correspond to realistic behaviour in the system being modelled. For example, if $Y$ represents physical size, $Y(t)<0$ is impossible. Negative growth is also possible if $Y(t)>\frac{\beta_{0}}{\beta_{1}}$.
Figure 1: Behaviour of numerical solutions at erroneous and true parameter locations showing considerable instability at the erroneous point. The true $Y(t_{j})$ values are those of the analytic solution with $\beta_{0}=10$ and $\beta_{1}=1$ that the model attempts to estimate, the erroneous values $\beta_{0}=49.3$ and $\beta_{1}=4.91$ can be found both in theoretical results and through empirical testing. (a) and (b) compare the estimated sizes for integer times between 0 and 4 to the analytic solution to show both the good and bad parameters give numerical values very close to the analytic ones. In (c) and (d) we see the sub-steps as well, which are connected up based on the numerical step they come from. The bad parameter estimates demonstrate considerable instability and even a negative value. (e) and (f) give the gradients at the numerical sub-steps and show why the estimated sizes are so unstable for the bad estimate as the gradient oscillates about 0.
@@
For this demonstration we use added error of $\mathcal{N}(0,0.1)$ and rounding to produce simulated data with measurement precision of 0.1, analogous to $1\text{\,}\mathrm{mm}$ measurement precision and standard deviation on observations in centimetres. The scale of measurement is chosen because the method in O’Brien et al., (2024) was developed for use in ecological size data where such structure shows up (Condit,, 1998). We want to demonstrate a realistic measurement process, but the bad numerical behaviour occurs even with a high level of precision, and much smaller error as the underlying cause is mathematical.
We are attempting to estimate the parameters $\beta_{0}$ and $\beta_{1}$ from simulated observations $y_{j}$ of the form in Equation (4). We estimate $\hat{Y}_{j}$ given the prior distribution
@@
and the longitudinal model in Equation (2), which encodes the auto-correlation in the data. For this demonstration we are looking at a single simulated individual. To make fitting the model easier we apply a translation by the mean observed size $\bar{y}$, which produces the implemented DE
@@
so the output estimates are consistent with Equation (1). We have default prior distributions for the parameters which are
@@
where the mean and standard deviation for $\beta_{1}$ are for the underlying log-transformed normal distribution. We chose a log-normal distribution as it enforces positive values to maintain our parameterisation of Equation (1). A prior sensitivity analysis for both centre and spread is provided in Supplement A and demonstrates that the bimodality is persistent under reasonable prior parameter choice. The minimal hierarchical set-up is based on a single individual structure such as used in O’Brien et al., (2024), but with a number chosen for the error standard deviation instead of fitting an error parameter.
@@ 2.1 Numerical integration theory
In this section we prove non-identifiability of the linear model parameters under a Runge-Kutta 4th order numerical method (RK4) in a theoretical context using the true values rather than observed data. We prove that, amongst the set of solutions for which the asymptotic size is $\alpha$, the linear ODE has multiple $\hat{\beta}_{1}$ parameters that will produce the same sizes over time for the RK4 method, then look at the specific case study where $\beta_{0}=10$ and $\beta_{1}=1$.
@@
In the case of Equation (1) the time parameter $t$ does not appear explicitly so our $k$ formulas do not depend on it. We will refer to where the $k$ parameters are estimated as ‘sub-steps’. Note that $t_{j}$ and $t_{j+1}=t_{j}+h$ are typically different to the times at which observations occur. The observation interval is an upper bound on the numerical step size $h$, but we can take multiple numerical steps between the observations.
In a chosen situation where $\beta_{0}=10$ and $\beta_{1}=1$ which produces a bimodal posterior, at the second (erroneous) parameter combination the numerical method is sufficiently unstable that forward projection produces $\hat{Y}(t_{j})$ values that match the analytic solution at the true parameters. For a visual comparison see Figure 1(a) and (b), which leverages the bad estimates ($\beta_{0}=49.3,\beta_{1}=4.91$) paired to our chosen parameters $\beta_{0}=10$ and $\beta_{1}=1$. The estimation algorithm will converge to any parameter combination that gives $\hat{Y}(t_{j})$ close to the observations provided. In these cases the algorithm finds a point far from the true value where size estimates match the observations extremely well because of numerical integration error. Non-identifiability here is not a problem with MCMC or another estimation method, it is a structural issue arising from the use of numerical integration in the longitudinal structure.
@@
The following provides proofs for non-identifiability in general for linear ODEs when using the RK4 numerical method. In the case of the linear ODE we can leverage an explicit form of the RK4 algorithm in Equation (5) to get a polynomial in $\beta_{0}$ and $\beta_{1}$. We prove non-identifiability for solutions that satisfy:
@@
For a first order linear ODE of the form
@@
$\hat{Y}(t_{j}+h)=Y(t_{j})+(\alpha-Y(t_{j}))\bigg(h\beta_{1}-\frac{h^{2}}{2}\beta_{1}^{2}+\frac{h^{3}}{6}\beta_{1}^{3}-\frac{h^{4}}{24}\beta_{1}^{4}\bigg)$ (6)
@@
[…]
@@
To extract the polynomial we substitute the $k$s into Equation 7 and do some re-organising of terms to get
@@
The linear DE with negative slope always converges to the asymptotic size $\beta_{0}/\beta_{1}=\alpha$, and in fact if we make the substitution $\alpha\beta_{1}=\beta_{0}$ we can reduce to a function of a single variable $\beta_{1}$ to get
@@
For a linear first order ODE of the form
@@
$0=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-1.$ (8)
@@
From the analytic solution in Equation (3) we know that
@@
which we can substitute into Equation (9) to get
@@
$=$ $\ (Y(t_{j})-\alpha)(e^{-\beta_{1}h}-1).$ (10)
When solving for $\hat{\beta}_{1}$ we re-arrange Equation (10) to get
@@
$=$ $\ (\alpha-Y(t_{j}))\big(f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1\big).$ (11)
As we assume $\alpha$ – the asymptotic size – is a constant, the only variable in Equation (11) is $\hat{\beta}_{1}$, so the solutions will be the same as those that solve
$f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1=0.$ (12)
@@
$Y(t_{j}+nh)-Y(t_{j})=f(\hat{\beta}_{1},h)\Bigg(\sum_{i=0}^{n-1}\alpha-Y(t_{j}+ih)\Bigg).$ (13)
As in the 1-step case we take the left hand side (LHS) of Equation (13) and substitute in the analytic solution $Y(t_{j}+nh)=\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}nh}$ to get
$\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}nh}-Y(t_{j})=(Y(t_{j})-\alpha)(e^{-\beta_{1}nh}-1),$ (14)
@@
$e^{-\beta_{1}nh}-1=\Bigg(\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg)(e^{-\beta_{1}h}-1),$ (15)
@@
Substituting Equation (15) into Equation (14) gives
$\text{LHS}=(Y(t_{j})-\alpha)\Bigg(\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg)(e^{-\beta_{1}h}-1).$ (16)
Let us return to the right side of Equation (13), and look at the argument within the sum. We substitute in the analytic solution to give
@@
$\text{RHS}=f(\hat{\beta}_{1},h)(\alpha-Y(t_{j}))\Bigg(\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg).$ (17)
Recombining Equations (16) and (17), we subtract Equation (16) from both sides to give
@@
$=$ $\ (\alpha-Y(t_{j}))\Bigg(\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg)(f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1).$ (18)
Equation (18) is 0 when
$0=f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-1,$ (19)
which is the same as Equation (12) from the 1-step case. As such, $\hat{\beta}_{1}$ solutions will be the same for all $n$. Equation (19) does not depend on $t_{j}$, so as long as $Y(t_{j})\neq\alpha$ the solutions are independent of $Y(t_{j})$ as well. ∎
@@
For a linear first order ODE of the form
@@
$0=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-1$
@@
For some parameter values Equation (8) may have a combination of real and complex roots, but as there must be at least one solution which approximates with the true (real) parameter $\beta_{1}$, we can exclude the case of all imaginary roots. If the discriminant of the polynomial in $\hat{\beta}_{1}$ is 0 there are some special cases where one real root (of some multiplicity) exists, that can be determined with further calculations when $h$ and $\beta_{1}$ are known. For example, for a given $h$ there will be one $\beta_{1}$ such that there is a single real root of multiplicity four given by
@@
and $\hat{\beta}_{1}$ can be determined through calculation using the discriminant when $h$ is chosen (Rees,, 1922). ∎
@@
For a linear first order ODE of the form
@@
As the true value $\hat{\beta_{1}}=\beta_{1}$ does not perfectly match the numeric solution to the analytic one, the estimate closest to $\beta_{1}$ that does solve Equation (8) will not be $\hat{\beta_{1}}=\beta_{1}$. ∎